OpenStudy (anonymous):

Find cot θ of csc θ = (sqrt 5)/2 and tan θ > 0.

zepdrix (zepdrix):

So like the last problem, let's figure out which quadrant we're in.

zepdrix (zepdrix):

Tangent is positive again, so we must be in the 1st or 3rd.

OpenStudy (anonymous):

1st

zepdrix (zepdrix):

\[\large \csc\theta \qquad=\qquad \frac{1}{\sin\theta} \qquad=\qquad \frac{\sqrt5}{2}\] First? Ok cool.

OpenStudy (anonymous):

yes i got that far!!

zepdrix (zepdrix):

\[\large \sin\theta=\frac{2}{\sqrt5}=\frac{opposite}{hypotenuse}\] Understand what I did with our fraction there?

OpenStudy (anonymous):

yes you got rid of the 1 on the numerator and had to flip the fraction to do so.

zepdrix (zepdrix):

Yah some flippin. Ok so let's draw our triangle.

zepdrix (zepdrix):



OpenStudy (anonymous):

okay

OpenStudy (anonymous):

opp=2 hyp=sqrt5

zepdrix (zepdrix):



zepdrix (zepdrix):

Ok, cool. So what's the missing side length? :D

OpenStudy (anonymous):

1

OpenStudy (anonymous):

so the answer is 1/2

OpenStudy (anonymous):

thank you!

zepdrix (zepdrix):

Yah that sounds right, good job.